OpenStudy (anonymous):

use the distributive property to simplify the expression -7(8x-3)+x

OpenStudy (anonymous):

-7(8x)= -56x -7(-3)= 21 x= x -56x + x + 21 = -55x +21

OpenStudy (turingtest):

@karinathabest Ignore the answer above, it makes no sense. distribute the -7 to the parentheses: -56x+21+x combine like terms (put the x's together with addition) -55x+21

OpenStudy (anonymous):

57x+21

OpenStudy (anonymous):

oops didnt distribute the negative sign 55x+21

OpenStudy (turingtest):

or that one...

OpenStudy (anonymous):

still missed it -55x+21

OpenStudy (turingtest):

there ya go, now we all agree. But you should explain your answer kretschy